OpenStudy (anonymous):

slope of the line containing the points (5, 9) and (-6, 1)

OpenStudy (zepp):

Use the point-slope form: \((y_1-y)=m(x_1-x)\)

OpenStudy (zepp):

Find the slope by using \(\large m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (zepp):

Can you find the slope?

OpenStudy (anonymous):

I Can Try

OpenStudy (zepp):

Good :)

OpenStudy (anonymous):

I Got 3/5 ?

OpenStudy (zepp):

This wrong :) Let me show you

OpenStudy (anonymous):

K

OpenStudy (zepp):

\(\large m=\frac{y_2-y_1}{x_2-x1}=\frac{1-9}{-6-5}=\frac{-8}{-11}=\frac{8}{11}\) This is what you should get

OpenStudy (anonymous):

Ok I messed up at -6 - 5 thanks

OpenStudy (zepp):

Now you got your slope, you can use the slope-intercept form to find the equation for the line passing through these two points! \(\large y=mx+b\\y=\frac{8}{11}x+b\) Can you take it from here?

OpenStudy (zepp):

No worries :)